OpenStudy (anonymous):

solve for B, show work -(3b - 15) = 6 ( 2b + 5 )

OpenStudy (aripotta):

first, do the distributive property. what do you get after doing that?

OpenStudy (aripotta):

do you know how to distribute?

OpenStudy (anonymous):

-3b - ( - 15) = 12b + 35

OpenStudy (aripotta):

6 x 5 = 30, not 35 :l and -3b - (-15) can be simplified to -3b + 15. so our new equation would be -3b + 15 = 12b + 30

OpenStudy (aripotta):

do you know what to do from there?

OpenStudy (anonymous):

my bad, you can simplify -3b - (-15) o.O how?

OpenStudy (aripotta):

yea, when you see two - like that, it means you're adding

OpenStudy (anonymous):

:D

OpenStudy (aripotta):

so -3b + 15

OpenStudy (anonymous):

then -3b - 12b = 30 + (-15)

OpenStudy (aripotta):

hm, that's not the way i would go about it, but yea, that works lol. now just simplify that so you can go to the next step

OpenStudy (anonymous):

okay, how would you do it though?

OpenStudy (aripotta):

oh, i would add 3b and subtract 30, but you'll get the same answer either way i think

OpenStudy (anonymous):

eh, not sure. -15b = 15 b = -1 ?

OpenStudy (aripotta):

yea, that's right

OpenStudy (anonymous):

cool, thanks

OpenStudy (aripotta):

no problem :)